# Graph Chirp Signal and Graph Fractional Vertex-Frequency Energy Distribution

### v1

@@ VII Numerical experiments
In this section, we conduct four numerical experiments to detect chirp signals from Examples 1–-4 using the proposed graph fractional vertex-frequency energy distribution domain filtering method (GFED-F), and apply this method to several real-world datasets, thereby validating the correctness of the theory.
@@ VII-A Detection of graph chirp signals
Example 3: For the signal $\mathbf{x}_{1}$ defined in Example 1, corrupted by complex circular Gaussian noise with a standard deviation of $\sigma=0.3$, the fractional order parameter is selected as the chirp rate of $\mathbf{x}_{1}$, with $a=0.7$.
Example 4: For the signal $\mathbf{x}_{2}$ defined in Example 1, corrupted by complex circular Gaussian noise with a standard deviation of $\sigma=0.3$, the fractional order parameter is selected as the chirp rate of $\mathbf{x}_{2}$, with $a=0.7$.
Example 5: For the signal $\mathbf{x}_{3}$ defined in Example 1, corrupted by complex circular Gaussian noise with a standard deviation of $\sigma=0.4$, the fractional order parameter is selected as the chirp rate of $\mathbf{x}_{3}$, with $a=0.6$.
Example 6: For the signal $\mathbf{x}_{4}$ defined in Example 1, corrupted by complex circular Gaussian noise with a standard deviation of $\sigma=0.4$, the fractional order parameter is selected as the chirp rate of $\mathbf{x}_{4}$, with $a=0.6$.
Figs. 811 plot both the GEDs and GFEDs of the noisy signals in Examples 3–6, along with those of the denoised signals after filtering. As observed, the GEDs and GFEDs of the noisy signals are significantly affected by noise, making it difficult to extract meaningful signal information. However, by applying the filtering method we proposed, effective denoising can be achieved in the vertex-fractional-frequency domain, improving the detection of chirp signals. Furthermore, detection in vertex-frequency domain is less effective due to the influence of spurious frequency components, leading to analysis errors. In contrast, the vertex-fractional-frequency domain enhances signal detection performance by providing more accurate vertex and frequency information, reducing such errors and allowing for better signal representation.
Figure 8: Detection of graph chirp signals through filtering: Case of $\mathbf{x}_{1}$. (a) The GED of the noisy signal. (b) The GFED (a=0.7) of the noisy signal. (c) The GED of the filtered signal. (b) The GFED (a=0.7) of the filtered signal.
Figure 9: Detection of graph chirp signals through filtering: Case of $\mathbf{x}_{2}$. (a) The GED of the noisy signal. (b) The GFED (a=0.7) of the noisy signal. (c) The GED of the filtered signal. (b) The GFED (a=0.7) of the filtered signal.
Figure 10: Detection of graph chirp signals through filtering: Case of $\mathbf{x}_{3}$. (a) The GED of the noisy signal. (b) The GFED (a=0.6) of the noisy signal. (c) The GED of the filtered signal. (b) The GFED (a=0.6) of the filtered signal.
Figure 11: Detection of graph chirp signals through filtering: Case of $\mathbf{x}_{4}$. (a) The GED of the noisy signal. (b) The GFED (a=0.6) of the noisy signal. (c) The GED of the filtered signal. (b) The GFED (a=0.6) of the filtered signal.
VII-B Denoise of real-data
We also conduct experiments on fiver real-world datasets: sea surface temperature (SST), particulate matter 2.5 (PM-25), the thickness data on the dendritic tree, the traffic volume data for Toronto, and the Minnesota road data. The first two datasets are discussed in [42], while the latter three datasets can be accessed in [43].
@@ VII-B Denoise of real-data
For the SST and PM-25 datasets, we use $k$-NN graph as in [42, 44] with $k=\{2,5,7\}$. Specifically, we use the SST data from months $T=\{50,120,270\}$, with zero-mean white Gaussian noise at noise levels $\sigma=\{15,45,60\}$. For the PM-25 data, we use data from days $T=\{50,120,270\}$, with zero-mean white Gaussian noise at noise levels $\sigma=\{15,25,35\}$. We compare the performance of the proposed GFED-F method with that of the optimal GFRT domain filtering method (OGFRFT-F) [45] at different fractional orders. Fig. 12 presents the MSE and SNR line plots of the two filtering methods under $k=5$. As can be seen, the filtering performance of both methods changes with the fractional-order parameter $a$. The GFED-F generally outperforms the OGFRFT-F for most fractional-order values. In fact, for several fractional orders, GFED-F achieves better results than the optimal filtering performance of OGFRFT-F.
Additionally, Table I provides the MSE and SNR values for both methods under the optimal parameters under $k=\{2,5,7\}$. In general, the GFED-F outperforms the OGFRFT-F in most scenarios. There are a few instances where OGFRFT-F shows slightly better performance. Specifically, for the SST dataset at $T=50$ and $T=120$ under the $7$-NN graph with a noise level of $\sigma=40$, the MSE values for GFED-F are higher than those for OGFRFT-F. Additionally, for the PM–25 dataset at $T=50$ under $2$-NN graph with a noise level of $\sigma=15$, the MSE and SNR values for GFED-F are slightly worse than those for OGFRFT-F. In all other cases, GFED-F consistently provides better performance, highlighting its superiority over OGFRFT-F.
$\sigma=15$
$T=50$, MSE Refer to caption
$T=120$, MSE Refer to caption
$T=270$, MSE Refer to caption
$T=50$, SNR Refer to caption
$T=120$, SNR Refer to caption
$T=270$, SNR Refer to caption
$\sigma=40$
$\sigma=65$
(a) SST data
$\sigma=15$
$\sigma=25$
$\sigma=35$
(b) PM-25 data
Figure 12: Comparison of MSE and SNR line plots for OGFRFT-F and GFED-F at different fractional orders using the $5$-NN graph. (a) SST data with noise levels $\sigma={15,45,60}$. (b) PM–25 data with noise levels $\sigma={15,25,35}$.
TABLE I: Comparison of MSE and SNR values of SST and PM–25 datasets under optimal parameters of OGFRFT-F and GFED-F.
@@
$7$–NN
MSE OGFRFT-F $28.3506$ ($a=1$) 48.5869 ($a=0.8$) $55.8849$ ($a=0.8$) $25.8131$ ($a=1$) 45.7667 ($a=0.8$) $53.2457$ ($a=0.8$) $23.4273$ ($a=1$) $49.7531$ ($a=0.8$) $58.1912$ ($a=0.8$)
GFED-F 28.0820 ($a=0.9$) $52.4520$ ($a=0.8$) 52.1758 ($a=0.7$) 25.3465 ($a=0.9$) $47.1676$ ($a=0.8$) 48.8752 ($a=0.7$) 22.6316 ($a=1.1$) 49.1505 ($a=0.8$) 55.5626 ($a=0.7$)
SNR OGFRFT-F $12.1921$ ($a=1$) $9.0967$ ($a=0.9$) $7.9153$ ($a=0.9$) $12.2400$ ($a=1$) $9.2752$ ($a=0.9$) $7.8787$ ($a=0.9$) $12.5604$ ($a=1$) $9.1663$ ($a=0.9$) $7.5997$ ($a=0.9$)
@@
$2$–NN
MSE OGFRFT-F 53.6247 ($a=0.1$) $72.5014$ ($a=0.1$) $86.6515$ ($a=0.1$) $30.8939$ ($a=0.3$) $39.0386$ ($a=0.4$) $47.2836$ ($a=0.5$) $14.6532$ ($a=0.9$) $15.7741$ ($a=1$) $16.2767$ ($a=1$)
GFED-F $56.2765$ ($a=0.3$) 64.5222 ($a=0.2$) 73.4911 ($a=0.1$) 29.2637 ($a=0.4$) 33.8578 ($a=0.4$) 37.8861 ($a=0.5$) 13.0029 ($a=0.3$) 13.2016 ($a=0.2$) 13.9401 ($a=0.2$)
SNR OGFRFT-F 5.4916 ($a=0.9$) $3.7791$ ($a=0.9$) $2.7184$ ($a=0.9$) $4.1291$ ($a=2$) $2.7812$ ($a=2$) $2.2033$ ($a=0.9$) $5.0868$ ($a=0.9$) $4.3523$ ($a=1$) $4.1595$ ($a=1$)
GFED-F $5.3348$ ($a=0.3$) 4.7409 ($a=0.2$) 4.1757 ($a=0.1$) 5.8716 ($a=0.4$) 5.2383 ($a=0.4$) 4.7501 ($a=0.5$) 5.6885 ($a=0.3$) 5.6227 ($a=0.2$) 5.3863 ($a=0.2$)
$5$–NN
MSE OGFRFT-F $56.2678$ ($a=1.6$) $69.4473$ ($a=0.6$) $76.3163$ ($a=0.6$) $32.1141$ ($a=0.8$) $38.1410$ ($a=0.8$) $40.8834$ ($a=0.8$) $14.2646$ ($a=2$) $16.5298$ ($a=0.7$) $15.7226$ ($a=0.9$)
@@
For the thickness data on the dendritic tree, the traffic volume data from Toronto, and the Minnesota road data, we use the same noise levels as in [46], namely $8\mathrm{dB}$, $7\mathrm{dB}$, and $5\mathrm{dB}$. Fig. 13 presents the SNR plots comparing the filtering method proposed in this paper with the GED-based Wiener filtering method (GED-WF) [46].
Figure 13: Comparison of GED-WF and GFED-F. (a) Dendritic tree. (b) Toronto traffic volume. (c) Minnesota road.
@@
VIII Conclusion

### v2

@@ VII Numerical experiments
In this section, we conduct numerical experiments to detect chirp signals from Examples 1–-2 using the proposed GFED-F, and apply this method to several real-world datasets, thereby validating the correctness of the theory.
@@ VII-A Detection of graph chirp signals
Example 3: For the signal $\mathbf{x}_{1}$ defined in Example 1, corrupted by complex circular Gaussian noise with a standard deviation of $\sigma=0.3$, the fractional order parameter is selected as the chirp rate of $\mathbf{x}_{1}$, i.e. $a=0.7$.
Example 4: For the signal $\mathbf{x}_{2}$ defined in Example 2, corrupted by complex circular Gaussian noise with a standard deviation of $\sigma=0.4$, the fractional order parameter is selected as the chirp rate of $\mathbf{x}_{2}$, i.e. $a=0.6$.
Figs. 78 present the GEDs and GFEDs of the noisy signals in Examples 1–2, as well as those of the denoised versions obtained using the proposed filtering method. As observed, both representations of the noisy signals are significantly affected by noise, which obscures the vertex-–frequency structure and hinders the identification of key signal features. However, the vertex–fractional-frequency domain clearly exhibits enhanced robustness to noise after filtering, improving the detection of chirp signals. In contrast, the filtered GEDs still suffer from spurious spectral activations, which lead to analysis errors in the vertex–-frequency domain. As shown in Figs. 7LABEL:sub@7c and 8LABEL:sub@8c, interference remains around vertex indices 30-–40 and frequency indices 50-–60. The corresponding Shannon entropy values of the filtered GEDs and GFEDs are listed in Table IV. The GFEDs consistently exhibit lower entropy values. These results confirm the effectiveness of the proposed filtering strategy in enhancing vertex–fractional-frequency representations for graph chirp signal detection.
TABLE IV: Shannon entropy values of the filtered GEDs and GFEDWs.
Signal Distribution Shannon entropy
$\mathbf{x}_{1}$ GED 45.042045.042045.0420
GFED ($a=0.7$) 44.4988
$\mathbf{x}_{2}$ GED 80.432980.432980.4329
GFED ($a=0.6$) 51.5346
Figure 7: Detection of graph chirp signals through filtering: Case of $\mathbf{x}_{1}$. (a) The GED of the noisy signal. (b) The GFED (a=0.7) of the noisy signal. (c) The filtered GED. (b) The filtered GFED (a=0.7).
Figure 8: Detection of graph chirp signals through filtering: Case of $\mathbf{x}_{2}$. (a) The GED of the noisy signal. (b) The GFED (a=0.6) of the noisy signal. (c) The GED of the filtered signal. (b) The GFED (a=0.6) of the filtered signal.
VII-B Denoising of real-data
We also conduct experiments on five real-world datasets: sea surface temperature (SST), particulate matter 2.5 (PM–25), the thickness data on the dendritic tree, the traffic volume data for Toronto, and the Minnesota road data. The first two datasets are discussed in [48], while the latter three datasets can be accessed in [49].
@@ VII-B Denoising of real-data
To evaluate the effectiveness of the proposed GFED-F, we compare it with both graph spectral filtering methods and graph neural network (GNN)–based filtering methods. Specifically, the spectral filtering methods include OGFRFT-F [46] and GED-WF [47], while the GNN–based methods involve three classical architectures: Chebyshev graph convolutional network (ChebNet) [50], graph attention network (GAT) [51], and graph convolutional network (GCN) [52].
Comparison with graph spectral filtering methods: For the SST and PM–25 datasets, we use $k$-NN graph as in [48, 53] with $k=\{2,5,7\}$. Specifically, we use the SST data from months $T=\{50,120,270\}$, with zero-mean white Gaussian noise at noise levels $\sigma=\{15,45,60\}$. For the PM–25 data, we use data from days $T=\{50,120,270\}$, with zero-mean white Gaussian noise at noise levels $\sigma=\{15,25,35\}$.
Fig. 9 presents the MSE and SNR line plots of the six filtering methods including OGFRFT-F, GED-WF, three GNN-based approaches (ChebNet, GAT, and GCN), and the proposed GFED-F. Figs. 9LABEL:sub@Fig9a and 9LABEL:sub@Fig9b depict the MSE and SNR results, respectively, for the SST dataset at $T=270$ with noise level $\sigma=40$. Figs. 9LABEL:sub@Fig9c and 9LABEL:sub@Fig9d depict the MSE and SNR results, respectively, for the PM–25 dataset at $T=50$ with noise level $\sigma=25$. The GFED-F consistently achieves strong performance across several fractional orders, and in many cases, it exceeds the best-case performance of OGFRFT-F. Notably, for certain fractional orders, GFED-F also outperforms GED-WF as well as all three GNN-based methods, further underscoring the effectiveness of the proposed framework for graph signal denoising.
Table V presents the MSE and SNR values for three spectral filtering methods: GED-WF, OGFRFT-F (optimized), and GFED-F (optimized), evaluated on the SST and PM–25 datasets under different graph connectivities ($k=2,5,7$) and noise levels. The results for OGFRFT-F and GFED-F are reported using their respective optimal fractional order parameters. Overall, GFED-F achieves the best performance in most settings, with lower MSE and higher SNR values compared to the other methods. In a few specific cases, such as the SST dataset at $T=50$ and $T=120$ under the 777-NN graph with $\sigma=40$, and the PM–25 dataset at $T=50$ under the 222-NN graph with $\sigma=15$, OGFRFT-F performs slightly better. Nevertheless, GFED-F consistently outperforms both OGFRFT-F and GED-WF in the majority of scenarios, demonstrating its robustness and effectiveness in graph signal denoising.
Figure 9: MSE and SNR comparison of six filtering methods on the SST dataset ($T=270$, $\sigma=40$) and the PM–25 dataset ($T=50$, $\sigma=25$). (a) MSE results on SST. (b) SNR results on SST. (c) MSE results on PM–25. (d) SNR results on PM–25.
TABLE V: Comparison of MSE and SNR values on the SST and PM–25 datasets using OGFRFT-F (Optimized), GED-WF and GFED-F (Optimized).
@@
777–NN
MSE OGFRFT-F 28.350628.350628.3506 ($a=1$) 48.5869 ($a=0.8$) 55.884955.884955.8849 ($a=0.8$) 25.813125.813125.8131 ($a=1$) 45.7667 ($a=0.8$) 53.245753.245753.2457 ($a=0.8$) 23.427323.427323.4273 ($a=1$) 49.753149.753149.7531 ($a=0.8$) 58.191258.191258.1912 ($a=0.8$)
GED-WF 42.082442.082442.0824 77.765677.765677.7656 119.2170119.2170119.2170 41.695341.695341.6953 78.367878.367878.3678 121.4248121.4248121.4248 46.726846.726846.7268 85.081285.081285.0812 134.8790134.8790134.8790
GFED-F 28.0820 ($a=0.9$) 52.452052.452052.4520 ($a=0.8$) 52.1758 ($a=0.7$) 25.3465 ($a=0.9$) 47.167647.167647.1676 ($a=0.8$) 48.8752 ($a=0.7$) 22.6316 ($a=1.1$) 49.1505 ($a=0.8$) 55.5626 ($a=0.7$)
SNR OGFRFT-F 12.192112.192112.1921 ($a=1$) 9.09679.09679.0967 ($a=0.9$) 7.91537.91537.9153 ($a=0.9$) 12.240012.240012.2400 ($a=1$) 9.27529.27529.2752 ($a=0.9$) 7.87877.87877.8787 ($a=0.9$) 12.560412.560412.5604 ($a=1$) 9.16639.16639.1663 ($a=0.9$) 7.59977.59977.5997 ($a=0.9$)
GED-WF 10.476810.476810.4768 7.80997.80997.8099 5.95445.95445.9544 10.157510.157510.1575 7.41707.41707.4170 5.51535.51535.5153 9.56209.56209.5620 6.95936.95936.9593 4.95824.95824.9582
@@
222–NN
MSE OGFRFT-F 53.6247 ($a=0.1$) 72.501472.501472.5014 ($a=0.1$) 86.651586.651586.6515 ($a=0.1$) 30.893930.893930.8939 ($a=0.3$) 39.038639.038639.0386 ($a=0.4$) 47.283647.283647.2836 ($a=0.5$) 14.653214.653214.6532 ($a=0.9$) 15.774115.774115.7741 ($a=1$) 16.276716.276716.2767 ($a=1$)
GED-WF 120.6260120.6260120.6260 211.4029211.4029211.4029 336.1629336.1629336.1629 65.881065.881065.8810 154.9327154.9327154.9327 260.6596260.6596260.6596 54.061654.061654.0616 121.8062121.8062121.8062 218.4704218.4704218.4704
GFED-F 56.276556.276556.2765 ($a=0.3$) 64.5222 ($a=0.2$) 73.4911 ($a=0.1$) 29.2637 ($a=0.4$) 33.8578 ($a=0.4$) 37.8861 ($a=0.5$) 13.0029 ($a=0.3$) 13.2016 ($a=0.2$) 13.9401 ($a=0.2$)
SNR OGFRFT-F 5.4916 ($a=0.9$) 3.77913.77913.7791 ($a=0.9$) 2.71842.71842.7184 ($a=0.9$) 4.12914.12914.1291 ($a=2$) 2.78122.78122.7812 ($a=2$) 2.20332.20332.2033 ($a=0.9$) 5.08685.08685.0868 ($a=0.9$) 4.35234.35234.3523 ($a=1$) 4.15954.15954.1595 ($a=1$)
GED-WF 2.02362.02362.0236 0.4131-0.4131- 0.4131 2.4275-2.4275- 2.4275 2.34732.34732.3473 1.3665-1.3665- 1.3665 3.6258-3.6258- 3.6258 0.5000-0.5000- 0.5000 4.0278-4.0278- 4.0278 6.5650-6.5650- 6.5650
GFED-F 5.33485.33485.3348 ($a=0.3$) 4.7409 ($a=0.2$) 4.1757 ($a=0.1$) 5.8716 ($a=0.4$) 5.2383 ($a=0.4$) 4.7501 ($a=0.5$) 5.6885 ($a=0.3$) 5.6227 ($a=0.2$) 5.3863 ($a=0.2$)
555–NN
MSE OGFRFT-F 56.267856.267856.2678 ($a=1.6$) 69.447369.447369.4473 ($a=0.6$) 76.316376.316376.3163 ($a=0.6$) 32.114132.114132.1141 ($a=0.8$) 38.141038.141038.1410 ($a=0.8$) 40.883440.883440.8834 ($a=0.8$) 14.264614.264614.2646 ($a=2$) 16.529816.529816.5298 ($a=0.7$) 15.722615.722615.7226 ($a=0.9$)
GED-WF 86.232086.232086.2320 97.612897.612897.6128 112.8108112.8108112.8108 40.419140.419140.4191 50.625050.625050.6250 64.951664.951664.9516 20.128320.128320.1283 32.255132.255132.2551 46.487946.487946.4879
@@
For the thickness data on the dendritic tree, the traffic volume data from Toronto, and the Minnesota road data, we use the same noise levels as in [47], namely $8\mathrm{dB}$, $7\mathrm{dB}$, and $5\mathrm{dB}$, respectively. Due to the relatively large scale of these datasets, the computational cost of OGFRFT-F is significantly higher. Therefore, we omit OGFRFT-F from this part of the comparison. Fig. 10 presents the MSE and SNR results comparing the proposed GFED-F with the GED-WF. It can be observed that for the dendritic tree data, our method achieves superior performance over GED-WF at the optimal fractional order. For the traffic volume and Minnesota road datasets, our method consistently outperforms GED-WF across multiple fractional orders. Fig. 11 further illustrates the denoising performance of GFED-F under the optimal fractional order through visualizations on the three datasets.
Figure 10: Comparison of GED-WF and GFED-F. (a) MSE results on Dendritic tree. (b) SNR results on Dendritic tree. (c) MSE results on Toronto traffic volume. (d) SNR results on Toronto traffic volume. (e) MSE results on Minnesota road. (f) SNR results on Minnesota road.
@@
[…]
TABLE VI: Comparison of MSE and SNR values on the SST and PM–25 datasets using ChebNet, GAT, GCN and GFED-F (Optimized).
SST $T=50$ $T=120$ $T=270$
$\sigma=15$ $\sigma=40$ $\sigma=15$ $\sigma=40$ $\sigma=15$ $\sigma=40$
MSE ChebNet 49.806049.806049.8060 69.098369.098369.0983 51.634451.634451.6344 69.747869.747869.7478 61.365961.365961.3659 83.339583.339583.3395
GAT 64.451664.451664.4516 68.520968.520968.5209 64.781164.781164.7811 65.859665.859665.8596 70.624570.624570.6245 75.979975.979975.9799
GCN 51.191451.191451.1914 59.255259.255259.2552 56.034556.034556.0345 60.885660.885660.8856 67.721667.721667.7216 74.443874.443874.4438
GFED-F 31.1213 56.9941 27.2796 51.9049 24.2385 49.7348
SNR ChebNet 9.74499.74499.7449 8.32318.32318.3231 9.22909.22909.2290 7.92317.92317.9231 8.37848.37848.3784 7.04927.04927.0492
GAT 8.62548.62548.6254 8.35958.35958.3595 8.24398.24398.2439 8.17228.17228.1722 7.76817.76817.7681 7.45077.45077.4507
GCN 9.62589.62589.6258 8.99058.99058.9905 8.87398.87398.8739 8.51338.51338.5133 7.95047.95047.9504 7.53947.53947.5394
GFED-F 11.7872 9.1595 12.0000 9.2063 12.4126 9.2911
PM–25 $T=50$ $T=120$ $T=270$
$\sigma=25$ $\sigma=35$ $\sigma=25$ $\sigma=35$ $\sigma=25$ $\sigma=35$
MSE ChebNet 113.1868113.1868113.1868 117.8153117.8153117.8153 45.132445.132445.1324 48.521948.521948.5219 22.252322.252322.2523 21.795321.795321.7953
GAT 113.0195113.0195113.0195 114.9533114.9533114.9533 47.282547.282547.2825 48.498248.498248.4982 18.766718.766718.7667 19.239719.239719.2397
GCN 119.8559119.8559119.8559 120.6042120.6042120.6042 50.738150.738150.7381 51.154851.154851.1548 20.857820.857820.8578 21.015521.015521.0155
GFED-F 59.4986 62.6643 35.6684 37.9269 14.0974 14.7062
SNR ChebNet 2.30012.30012.3001 2.12602.12602.1260 3.99003.99003.9900 3.67553.67553.6755 3.35523.35523.3552 3.44533.44533.4453
GAT 2.30652.30652.3065 2.23282.23282.2328 3.78793.78793.7879 3.67773.67773.6777 4.09504.09504.0950 3.98693.98693.9869
GCN 2.05142.05142.0514 2.02442.02442.0244 3.48163.48163.4816 3.44613.44613.4461 3.63623.63623.6362 3.60353.60353.6035
GFED-F 5.0930 4.8678 5.0121 4.7455 5.3375 4.8040
VIII Discussion
A key observation from our study is that graph chirp signals are inherently dependent on the underlying graph topology. Different graph structures naturally give rise to different graph chirp signals. While the underlying graph is often constructed using conventional methods such as the $k$-NN graph, recent works have proposed data–driven approaches that learn the graph structure by optimizing the graph representation matrix [54, 55, 56]. This insight leads to a novel perspective in which any real-world signal $\mathbf{x}$ can be interpreted as a graph chirp signal defined on a suitable graph structure. Given a signal $\mathbf{x}\in\mathbb{R}^{N}$, we first normalize it to obtain $\tilde{\mathbf{x}}$. Then, with a chosen graph chirp rate $a$ and initial graph frequency $k$, we extend $\tilde{\mathbf{x}}$ into a complete orthonormal basis of $\mathbb{R}^{N}$. This can be accomplished using various techniques such as Gram–Schmidt orthogonalization, Householder transformations, QR decomposition, or singular value decomposition. By inserting $\tilde{\mathbf{x}}$ as the $k$-th basis vector, we form the inverse GFRFT matrix
$(\mathbf{F}_{G}^{a})^{-1}=\left[\mathbf{u}_{1}^{a},\cdots,\mathbf{u}_{k-1}^{a},\tilde{\mathbf{x}},\mathbf{u}_{k+1}^{a},\cdots,\mathbf{u}_{N}^{a}\right].$ (52)
The matrix
$\mathbf{U}=\left((\mathbf{F}_{G}^{a})^{-1}\right)^{\frac{1}{a}}=\left[\mathbf{u}_{1}^{a},\cdots,\mathbf{u}_{k-1}^{a},\tilde{\mathbf{x}},\mathbf{u}_{k+1}^{a},\cdots,\mathbf{u}_{N}^{a}\right]^{\frac{1}{a}}$ (53)
can then be treated as the eigenvector matrix of a graph shift operator. By assigning a suitable diagonal eigenvalue matrix $\mathbf{\Lambda}$ based on the desired task objective (e.g., smoothing, denoising, compression), we can construct a graph shift operator $\mathbf{Z=U\Lambda U}^{-1}$. This approach provides a practical and adaptable way to build graph representations directly from data, making it broadly applicable to real-world signal processing problems.
Moreover, the proposed filtering framework can be extended to non–linear graph signal models. For example, it is worth exploring whether our approach can be adapted to frameworks such as the one proposed in [57]. Future work may investigate how to incorporate fractional-frequency representations into such non-linear settings, possibly by learning graph chirp rates in a model–aware manner or integrating them with advanced inference schemes.
IX Conclusion